OpenStudy (anonymous):

Simplify the given expression: (1 + 2i) • (5 − 3i) 2 −15i 3 + 2i 5 −6i 11 +7i

OpenStudy (one098):

18i ??

OpenStudy (anonymous):

\[\left( a+bi \right) \cdot \left( c+di \right)=\left( ac-bd \right)+\left( ad+bc \right)i\]

OpenStudy (anonymous):

11+ 7i?

OpenStudy (anonymous):

that's what i get